a. In a butter-packing plant, the quality of butter packed in a day using a certain type of machine is normally distributed with mean and variance 4. On a particular day, 12 packets of butter were taken at random from this production line and their masses, measured in grams, were:
9.5 11.2 9.9 10.9 10.1 10.9
9.5 10.6 11.1 9.8 10.2 11.0
Find a 97% confidence interval for the mean mass produced by this machine.
Explain the answer.

b. What sample size would be required to estimate the population mean for a large file of invoices of a multinational company to within RM0.50 with 95% confidence, given that the estimated standard deviation of the value of the invoices is RM6?
1
Expert's answer
2020-04-29T16:34:08-0400

"a)\hat{\mu}\approx 10.39\text{ --- point estimate of population mean}.\\
\sigma^2=4.\\
N=12.\\
\alpha=0.03.\\
z_{\alpha/2}=2.17.\\
\mu=\hat{\mu}\pm [z_{\alpha/2}\frac{\sigma}{\sqrt{N}}].\\
\mu=10.39\pm[2.17{\frac{2}{\sqrt{12}}}].\\
(9.14;11.64)\text{ --- our confidence interval}.\\
\text{We can say that this CI covers population mean with probability } 0.97.\\
b)\sigma=6.\\
\alpha=0.05.\\
z_{\alpha/2}=1.96.\\
z_{\alpha/2}\frac{\sigma}{\sqrt{N}}=0.5.\\
N=(\frac{z_{\alpha/2}\sigma}{0.5})^2.\\
N\approx 554."